I know it looks like it's two questions but according my book it is one. Please help me.
A function is defined by f(x) = 2x - 4
f (x + h)
f (3)

Answer by bucky(2189)

I'm not sure what your book means when it says this is one problem. But let's discuss how
to do this problem.
.
We are given:
.
f(x) = 2x - 4
.
Let's do the last one first. We are asked to find f(3). All this is telling us to do is
to go to the given equation, substitute 3 wherever you see x, and then simplify.
.
So to find f(3) we start with the given equation:
.
f(x) = 2x - 4
.
Then we substitute into that equation a 3 for every x and we get:
.
f(3) = 2*3 - 4
.
Then we simplify the right side by multiplying the 2 times 3 to get 6 which makes the equation:
.
f(3) = 6 - 4
.
and finally we subtract the 4 from the 6 to get:
.
f(3) = 2
.
That's the answer for f(3). To get f(x + h) we do the same thing, only this time we substitute
x + h for every x in the given equation. So start with the given equation:
.
f(x) = 2x - 4
.
Now replace every x in that equation with x + h and we get:
.
f(x + h) = 2*(x + h) - 4
.
Now multiply out the right side to get:
.
f(x + h) = 2x + 2h - 4
.
and this is as much as we can simplify the right side. So the answer is:
.
f(x + h) = 2x + 2h - 4
.
Hope this helps you to understand what the problem is asking you to do, and helps you
to understand what you do to solve problems such as this.
.